# LOGIC-LM++: Multi-Step Refinement for Symbolic Formulations

### v1

@@ Abstract
In this paper we examine the limitations of Large Language Models (LLMs) for complex reasoning tasks. While current approaches leverage formal languages as intermediate representation of reasoning problems, they struggle with generating intermediate formal specifications and refining these representations. To address these issues, this paper proposes Logic-LM++ Pan et al. (2023), an improvement on Logic-LM. It uses the ability of LLMs to do pairwise comparisons, allowing the evaluation of the refinements suggested by the LLM. The paper demonstrates that Logic-LM++ outperforms Logic-LM and LLM based techniques on natural language reasoning tasks on two datasets, 1 and AR-LSAT. Logic-LM++ show an average improvement of 13.5% on standard prompting, 11% on chain of thought prompting and 5% on Logic-LM.
@@
Shashank Kirtania, Priyanshu Gupta, Arjun Radhakrishna Microsoft
@@ 1 Introduction
While Logic-assisted LLM reasoning techniques are promising, we observe following problems in such systems: Firstly LLMs are poor at generating intermediate formal specifications. A few techniques try to counter this problem with a refinement loop Madaan et al. (2023a); Welleck et al. (2022); Shinn et al. (2023) to improve on the syntactical correctness of the symbolic formulation. Secondly, the LLMs are poor at refinement of the formal representations in Figure 1 we demonstrate to generate appropriate formulations for the problem, LLM takes an extra step of refinement. Even when the LLM generated symbolic formulation is almost correct, during refinement stage the LLM introduced a new error by incorrectly translating the statement "No young person teaches". Incorrect translation from Natural Language (NL) to intermediate formal specifications is a common problem we observe over the failing cases of refinement. Thirdly, we observe that refinements are not always linear- resolving an error with the symbolic formulation can take multiple steps of careful edits and evaluation. The formulations generated in refinement stage introduced the wrong interpretation of "No young person teaches" to "All young people teaches" with the backtracking agent we are able to revert back to previous formulation since it was semantically correct.
To address these challenges we propose Logic-LM++ that develops upon Logic-LM
1.
We leverage the ability of LLMs to do pairwise comparison Zheng et al. (2023a), this gives us an opportunity to evaluate the refinements suggested by the LLM and do a semantic check with respect to the problem statement to ensure the edits in the symbolic formulation improve formulation semantically.
2.
We improve on the refinement present in Logic-LM to give more context on the problem statement during refinement, this eliminates cases where recommended edits are appalling.
@@ 2.1 Reasoning with LLMs
Language model reasoning (LLM) commonly entails the deconstruction of complex questions into a sequence of intermediate steps, often referred to as chains, before reaching the ultimate solution. This technique is a reflection of methods such as Chain of Thought (CoT) prompting and its variations, as shown in various studies Wei et al. (2022); Kojima et al. (2022). These methodologies require the meticulous segmentation of a problem into a chain of smaller, manageable chunks. Each of these chunks represents a step in the reasoning process, guiding the model towards a comprehensive solution. The concept of the reflection loop, as explored in previous research Shinn et al. (2023); Madaan et al. (2023b), offers a means of refining the reasoning by identifying and eliminating any flaws that may be introduced by the LLM during a reasoning step. This process enhances the inherent capability of the LLM to self-correct, contributing to more accurate and reliable outcomes.
@@
Dataset GPT-3.5 Turbo GPT-4 Standard CoT Logic-LM Logic-LM++ Standard CoT Logic-LM Logic-LM++ 1 45.09 57.35 62.80 62.25 69.11 70.58 78.92 84.80 AR-LSAT 20.34 17.31 26.41 28.13 33.33 35.06 43.04 46.32
@@ 3.1 Technique
Result interpretation where the produced output is mapped to the right answer using regex parsing.
Logic-LM uses a refinement loop to fix errors in symbolic formulation at formulation and reasoning stages. However, Logic-LM still struggles to improve on logical representations, showing almost no improvement after multiple iterations. Authors attribute this to semantic limitations of the formulation. This is where Logic-LM++ comes in, it enhances the semantic representation of NL problem in the symbolic formulations.
@@ 3.2 Self-Refinement Agent
We build upon the self-refinement agent used in Logic-LM. In Figure 1 we show how the refinement prompt is designed. We build upon the self-refinement agent of Logic-LM by introducing the problem statement and the question to give LLM more context to repair the formulation, this enables us to remove the cases where the LLM generates refinement that hampers the structure of the formulation as shown in Figure 1, we show a sample of the refinement prompt that is sent to the LLM to use to come up with the right formulations. Originally Logic-LM uses a few-shot approach to refine the formulations, this brings a lot of extra irrelevant information about repairing the bug relevant to the current problem statement.
@@ 3.3 Backtracking Agent
LLMs has shown remarkable results in automated evaluating benchmarks Zheng et al. (2023b) and has shown high alignment with the human judgement Wei et al. (2024). We use this capability of LLMs to assess if the repair by self-refinement improves the alignment of the human intent with LLM generated formulations. This allows us to get rid of the updates that are not helpful in future iterations and only use those updates where the changes help the model to come to the right formulation.
@@ 4.1 Dataset
1 Han et al. (2022) is a challenging expert-written dataset for logical reasoning. The problems are aligned with real-world knowledge and use highly natural wordings, and the questions require complex first-order logic reasoning to solve. We use the entire 1 test set for evaluation with 204 examples. AR-LSAT Zhong et al. (2022) is a dataset that collects all analytical logic reasoning questions from the Law School Admission Test from 1991 to 2016. We use the test set which has 231 multiple-choice questions. AR-LSAT is particularly challenging, with state-of-the-art model’s performance slightly better than random guessing Liang et al. (2022); Ribeiro et al. (2023).
@@ 4.2 Principal Findings
| BT | FOLIO | AR-LSAT | |||
| exe rate | exe acc | exe rate | exe acc | ||
| GPT-3.5 turbo | - | 84.3 | 64.3 | 21.8 | 60.3 |
| + | 84.3 | 66.2 | 22.9 | 64.1 | |
| GPT-4 | - | 85.8 | 79.9 | 32.6 | 60.0 |
| + | 86.7 | 85.8 | 32.0 | 66.2 | |
@@
RQ2: Does refinement by LLM always positively affect the formulations? We evaluate the refinement with and without backtracking step in Figure 3. It is evident that as the number of runs increase the accuracy flatten for Logic-LM, the main reason behind that is the refined solution might be executable but is not a valid representation of the code itself. With the possibility of reverting back to the initial code in case the refined code is not semantically better the Logic-LM++ show consistently better results over multiple runs.
@@ 6 Conclusion
We propose Logic-LM++ which beats state of the art results on natural language reasoning tasks on two datasets. Logic-LM++ takes leverage of LLMs’ reasoning capabilities to show significant improvements in efficient use of logic solvers for reasoning, we demonstrate that LLMs show promising results at conducting comparison between symbolic formulations even in cases where generating symbolic formulations is a hard task for LLM.
@@
At present, Logic-LM++ faces constraints in its capacity to effectively capture the semantic intricacies inherent in reasoning tasks. This limitation notably complicates the evaluation process, especially when dealing with smaller LLMs like Rozière et al. (2023). The nuanced understanding required for accurate reasoning poses a significant challenge, particularly in contexts where the model’s semantic comprehension may be insufficient. As a consequence, the assessment of performance, particularly in comparison to more compact language models, becomes notably more complex. This limitation underscores the need for continued advancements in semantic understanding within language models to enhance their efficacy across diverse reasoning tasks.

### v2

@@ Abstract
In this paper we examine the limitations of Large Language Models (LLMs) for complex reasoning tasks. Although recent works have started to employ formal languages as an intermediate representation for reasoning tasks, they often face challenges in accurately generating and refining these formal specifications to ensure correctness. To address these issues, this paper proposes Logic-LM++, an improvement on Logic-LM Pan et al. (2023). It uses the ability of LLMs to do pairwise comparisons, allowing the evaluation of the refinements suggested by the LLM. The paper demonstrates that Logic-LM++ outperforms Logic-LM and other contemporary techniques across natural language reasoning tasks on three datasets, 1, ProofWriter and AR-LSAT, with an average improvement of 18.5% on standard prompting, 12.3% on chain of thought prompting and 5% on Logic-LM.
@@
Shashank Kirtania, Priyanshu Gupta, Arjun Radhakrishna Microsoft {t-skirtania, priyansgupta, arradha} @microsoft.com
@@ 1 Introduction
While logic-assisted LLM reasoning techniques are promising, we observe following problems in such systems: Firstly, LLMs are poor at generating intermediate formal specifications. A few techniques try to counter this problem with a refinement loop Madaan et al. (2023a); Welleck et al. (2022); Shinn et al. (2023) to improve upon the syntactical correctness of the symbolic formulation. Secondly, the LLMs are poor at repairing the formal representations with limited information with error information. For example, in Figure 1 the LLM initially generates a syntactically incorrect formulation. After a turn of refinement, while the LLM is able to generate a response that is syntactically correct, it introduces a semantic error in the formulation by incorrectly translating the statement "No young person teaches". These kind of incorrect translations from Natural Language (NL) to intermediate formal specifications is a common problem we observe over the failing cases of refinement. Thirdly, we observe that refinements are not always linear-resolving an error with the symbolic formulation can take multiple steps of careful edits and evaluation. The formulations generated in refinement stage in 1 introduced the wrong interpretation of "No young person teaches" to "All young people teaches".
To address these challenges we propose to add following measures in Logic-LM to enhance it’s capabilities resulting in improved variant Logic-LM++.
We leverage the ability of LLMs to do pairwise comparison Zheng et al. (2023a), this gives us an opportunity to evaluate the refinements suggested by the LLM and do a semantic check with respect to the problem statement to ensure if the edits in the symbolic formulation generated while refinement improve the formulation semantically not just syntactically.
We also improve on the refinement mechanism present in Logic-LM to give more context of the problem statement during refinement stage, this eliminates cases where recommended edits are appalling and do not improve the formulation significantly.
@@ 2.1 Reasoning with LLMs
Large Language model (LLM) - based reasoning techniques commonly entail the deconstruction of complex questions into a sequence of intermediate steps, often referred to as chains, before reaching the ultimate solution. This technique is a reflection of methods such as Chain of Thought (CoT) prompting and its variations, as shown in various studies Wei et al. (2022); Kojima et al. (2022). These methodologies require the meticulous segmentation of a problem into a chain of smaller, manageable chunks. Each of these chunks represents a step in the reasoning process, guiding the model towards a comprehensive solution. The concept of the reflection loop, as explored in previous research Shinn et al. (2023); Madaan et al. (2023b), offers a means of refining the reasoning by identifying and eliminating any flaws that may be introduced by the LLM during a reasoning step. This process enhances the inherent capability of the LLM to self-correct, contributing to more accurate and reliable outcomes.
@@
| Dataset | GPT-3.5 Turbo | GPT-4 | ||||||
| Standard | CoT | Logic-LM | Logic-LM++ | Standard | CoT | Logic-LM | Logic-LM++ | |
| 1 | 45.09 | 57.35 | 62.80 | 62.25 | 69.11 | 70.58 | 78.92 | 84.80 |
| ProofWriter | 35.50 | 49.17 | 58.33 | 58.83 | 52.67 | 68.11 | 79.66 | 79.66 |
| AR-LSAT | 20.34 | 17.31 | 26.41 | 28.13 | 33.33 | 35.06 | 43.04 | 46.32 |
@@ 3.1 Background
Result interpretation, where the produced output is mapped to the right answer using regex parsing.
Logic-LM uses a refinement loop to fix errors in symbolic formulation at formulation and reasoning stages. However, Logic-LM still struggles to improve on logical representations, showing almost no improvement after multiple iterations. Authors attribute this to semantic limitations of the formulation. To this end, Logic-LM++ aims to mitigate this limitation by improving the Logic-LM refinement loop.
@@ 3.2 Self-Refinement Agent
Logic-LM defines the notion of a Self-Refinement Agent to implement the refinement loop in the symbolic formulations in cases where the formulations did not yield a successful execution within the system. This agent is characterized by a refinement prompt 1 PG: Add more details in figure caption. In the original work, the refinement prompt constituted various few shot examples to act as exemplar for the model. While similar techniques have proven useful Madaan et al. (2023a); Shinn et al. (2023), we anecdotally observe that instead of helping the model it adds extra irrelevant information that distracts the model from fixing the issues relevant to the current formulation, consistent with similar studies in other domains Pan et al. (2023). To alleviate this, instead of adding few-shots, we add the problem statement and the question to the refinement prompt alongside instructions to self-reflect on the model’s failure to generate the right response. As we show later in Section 4, this structure helps better contextualize Shinn et al. (2023) the formulation to the self-reflection agent and help the system generate better refinements.
@@ 3.3 Backtracking Agent
LLMs has shown remarkable results in automated evaluation benchmarks Zheng et al. (2023b) and has shown high alignment with the human judgement Wei et al. (2024). We use this capability of LLMs to assess if the repaired formulation by self-refinement improves the alignment of the human intent with LLM generated formulations. This allows us to get rid of the updates that are not helpful in future iterations and only use those updates where the changes help the model to come to the right formulation.
@@ 4.1 Dataset
1 Han et al. (2022) is a challenging expert-written dataset for logical reasoning. The problems are aligned with real-world knowledge and use natural wordings, and the questions require complex first-order logic reasoning to solve. We use the 1 test set for evaluation with 204 examples.
AR-LSAT Zhong et al. (2022) is a dataset that collects all analytical logic reasoning questions from the Law School Admission Test from 1991 to 2016. We use the test set which has 231 multiple-choice questions. AR-LSAT is particularly challenging, with state-of-the-art model’s performance slightly better than random guessing Liang et al. (2022); Ribeiro et al. (2023).
ProofWriter Tafjord et al. (2021) is another popular dataset on deductive logical reasoning. Since the problems are in more natual language like setting it makes semantic evaluation very relevant in the problem set. Logic-LM use the open-world assumption (OWA) subset in which each example is a (problem, goal) pair and the label is one of PROVED, DISPROVED, UNKNOWN. Logic-LM evaluate the pipeline with the hardest section of ProofWriter which contain total of 600 randomly sampled five step multi-hop reasoning questions.
@@ 4.2 Principal Findings
| GPT-3.5 turbo | GPT-4 | ||||
| BT | - | + | - | + | |
| FOLIO | 84.3 | 84.3 | 85.8 | 86.7 | |
| 64.3 | 66.2 | 79.9 | 85.8 | ||
| ProofWriter | 95.6 | 95.6 | 99.0 | 99.0 | |
| 74.1 | 77.2 | 79.6 | 79.6 | ||
| AR-LSAT | 21.8 | 22.9 | 32.6 | 32.0 | |
| 60.3 | 64.1 | 60.0 | 66.2 | ||
@@
RQ2: Does refinement by LLM always positively affect the formulations?
In Figure 3, we evaluate the refinement process with and without backtracking. Logic-LM’s accuracy plateaus with more runs because refined solutions may not represent the intended code. The author’s also discuss this as a known limitation of the refinement process in the refinement loop they proposed. Backtracking, which reverts to the initial code if no semantic improvement is found, allows Logic-LM++ to perform consistently better by continually reassessing and correcting refinements for more reliable results.
Figure 4 shows that the backtracking agent significantly improves results in the second round within the 1 dataset, with a similar impact in later rounds. This indicates that backtracking is most effective early on since the generated refinement can also degrade the performance of the formulations, enabling Logic-LM++ to achieve substantial better and iterative improvements over time.
4.3 Error Analysis
Even though Logic-LM++ shows impressive improvements over standard refinement techniques, it still lacks behind in the cases where the first set of formulation generated is completely different from the ground truth formulation. On analyzing the failure cases in Logic-LM we note that the current pipeline relies a lot on fixing the bugs with current formulation without losing on semantic understanding, however in cases where the generating semantically correct formulations is hard the technique is contingent to initial formulations generated.
@@ 6 Conclusion
We propose Logic-LM++ which beats state of the art results on natural language reasoning tasks on three datasets. Logic-LM++ takes leverage of LLMs’ reasoning capabilities to show significant improvements in efficient use of logic solvers for reasoning, we demonstrate that LLMs show promising results at conducting comparison between symbolic formulations even in cases where generating symbolic formulations is a hard task for LLM.
@@
At present, Logic-LM++ faces constraints in its capacity to effectively capture the semantic intricacies in reasoning tasks. This limitation notably complicates the evaluation process, especially when dealing with smaller LLMs like Rozière et al. (2023). The understanding required for accurate reasoning poses a significant challenge, particularly in contexts where the model’s semantic comprehension may be insufficient. Due to this the assessment of performance becomes notably more complex. This limitation underscores the need for continued advancements in semantic understanding within LLMs to enhance their efficacy across reasoning tasks.